find the length of a rectangle if the perimeter is 72 yd and the length is three times the width
Answer by ChrisDS(7)

Perimeter=2l+2w
p=72
2l+2w=72
l=3w
2(3w)+2w=72
6w+2w=72
8w=72
w=9
l=3w
l=3(9)
l=27
Length=27yd
Width=9yd